40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 29 17:46:43 2018, gautam, Update, General, Some to-dos 
    Reply  Fri Nov 30 15:53:52 2018, Jon, Omnistructure, General, N2 pressure gauge fix N2_transducer_datasheet.pdf
       Reply  Tue Dec 4 18:24:05 2018, gautam, Omnistructure, General, N2 line disconnected IMG_7195.JPG
          Reply  Thu Dec 6 17:33:33 2018, Jon, Omnistructure, General, N2 line disconnected 
Message ID: 14324     Entry time: Thu Nov 29 17:46:43 2018     Reply to this: 14325
Author: gautam 
Type: Update 
Category: General 
Subject: Some to-dos 

[koji, gautam, jon, steve]

  • We suspect analog voltage from N2 pressure gauge is connected to interfacing Omega controller with the 'wrong' polarity (i.e pressure is rising over ~4 days and then rapidly falling instead of the other way around). This should be fixed.
  • N2 check script logic doesn't seem robust. Indeed, it has not been sending out warning emails (threshold is set to 60 psi, it has certainly gone below this threshold even with the "wrong" polarity pressure gauge hookup). Probably the 40m list is rejecting the email because controls isn't a part of the 40m group.
  • Old frames have to be re-integrated from JETSTOR to the new FB in order to have long timescale lookback.
  • N2 cylinder pressure gauges (at the cylinder end) need a power supply - @ Steve, has this been purchased? If not, perhaps @ Chub can order it.
  • MEDM vacuum screen should be updated to have gate valves be a different color to the spring-loaded valves. Manual valve between TP1 and V1 should also be added.
  • P2, P3 and P4 aren't returning sensible values (they should all be reading ~760torr as is P1). @ Steve, any idea if these gauges are broken?
  • Hornet gauges (CC and Pirani) should be hooked up to the new vacuum system.
  • add slow channels of   foreline pressures of TP2 & 3   and    C1:Vac-IG1_status_pressure
ELOG V3.1.3-